If 45.0 mL of 1.78M ethos like is combusted with excess oxygen to produce 1.00 g of water, what is the percent yield of this reaction?
1
Expert's answer
2020-01-27T05:14:23-0500

If 45.0 mL of 1.78M ethos like is combusted with excess oxygen to produce 1.00 g of water, what is the percent yield of this reaction?

V = 45.0 mL

C = 1.78 M

m (H2O) = 1.00 g

Solution:

С2H6О + 3O2 -> 2CO2 + 3H2O

n ( H2O) = m(H2O)/ M(H2O) =1.00g /18 g/mol = 0.06 mol

n teoretical ( C2H6O ) = V*C = 0.045L* 1.78 M = 0.08 mol

n practical (C2H6O ) = n (H2O)/3 = 0.06 mol/3 = 0.02 mol

n practical/n teoretical = 0.02 mol/ 0.08 mol *100% = 25%

Answer: 25%